solve the system of equations using the substitution method
x+z=8
y-z=5
x-y=9

Answer by Theo(13342)

your 3 equations are:
x+z=8 (first equation)
y-z=5 (second equation)
x-y=9 (third equation)
from the first equation you get x = 8-z
from the second equation you get y = z+5
substitute for x and y in the third equation to get:
8-z - (z+5) = 9
simplify to get:
8-z-z-5 = 9
combine like terms to get:
3-2z = 9
subtract 3 from both sides of the equation to get:
-2z = 6
divide both sides of the equation by -2 to get:
z = -3
substitute for z in first equation to get:
x = 11
substitute for z in the second equation to get:
y = 2
your solution set is:
x = 11
y = 2
z = -3
substitute in all 3 original equations to get:
x+z=8 becomes 11-3 = 8 which becomes 8 = 8
y-z=5 becomes 2+3 = 5 which becomes 5 = 5
x-y=9 becomes 11-2 = 9 which becomes 9 = 9
solutions are confirmed as good.